A drug tester claims that a drug cures a rare skin disease 68% of the time. The claim is checked by testing the drug on 100 patients. If at least 62 patients are​ cured, the claim will be accepted. Find the probability that the claim will be rejected assuming that the​ manufacturer's claim is true. Use the normal distribution to approximate the binomial distribution if possible.

The probability is __________(round to four decimal places as needed.)
Answer #1

Mean = np = 100 * 0.68 = 68

Standard deviation = sqrt (np(1-p) )

= sqrt (100 * 0.68 * 0.32)

= 4.6648

Using normal approximation,

P(X >= 62) = P(X > (61.5 - 68) / 4.6648) (Using continuity correction )

= P(Z > -1.39)

= P(Z < 1.39)

= 0.9177